Step by step
The question
Question: x can be 1, 2, 3, with probabilities 0.5, a missing one, 0.2: find the missing probability.
x: 1, 2, 3; P(x): 0.5, ?, 0.2
1 All P(x) add to 1
x always takes one of its values.
So all the probabilities add up to 1.
a probability table always adds up to 1
0.5 plus P of 2 plus 0.2 is 1.
0.5 + P(2) + 0.2 = 1
The known ones add up to 0.7.
0.7 + P(2) = 1
To remove the plus 0.7, we take away 0.7 on both sides.
1 minus 0.7 is 0.3.
P(2) = 1 − 0.7 = 0.3
Careful!
The missing probability is what is left over from 1.
Careful: the missing P is what is left of 1
2 Check
Let's check: 0.5 plus 0.3 plus 0.2 is 1.
So it is right!
0.5 + 0.3 + 0.2 = 1
Let's look back.
First: all P(x) add to 1.
Last: check.
Answer
Careful: the missing P is what is left of 1